OpenStudy (anonymous):

Two physics students are on a balcony 19.6 m above the street. One student throws a ball vertically downward at 14.7 m/s; at the same instant, the other student throws a ball vertically upward at the same speed. The second balljust misses the balcony on the way down. (a) What is the difference in the two balls' time in the air? (b)What is the velocity of each ball as it strikes the ground? (c) How far apart are the balls 0.8 s after they are thrown?

OpenStudy (anonymous):

you need to understand position and velocity equation for this problem \[x(t)=\frac{1}{2}at^2+v_ot+x_o\] \[v(t)=at+v_o\] now for the general constants acceleration (a) is equal to acceleration of gravity(g) =-9.81 m/s^2 the initial position x_o =19.6m

OpenStudy (anonymous):

now 2 cases if you throw the ball upwards, then v_o= 14.7 if you throw the ball downwards, then v_o=-14.7 to determine a) substitute the values into the position equation and determine t when x(t)=0 for both cases x(t)=0 means when the balls reach the ground or hit the ground and the corresponding t value would be the time it reaches the ground now the difference of the 2 times would be your answer for a

OpenStudy (anonymous):

now for b) once you have the 2 t values for once they reach the ground substitute the corresponding values into the velocity equation given above to determine the velocity at which it hits the ground for each of the time values basically solve v(t) for both t values

OpenStudy (anonymous):

c) for the 2 cases substitute the respective values into the position equation, then determine x(.8) x(.8) is the y position of the ball when the time is equal to .8 seconds then find the difference between the 2 position values

OpenStudy (anonymous):

@K0GA !! :D

OpenStudy (anonymous):

yea :)

OpenStudy (anonymous):

perfect completeidiot

OpenStudy (anonymous):

i can't understand what you want from a) tell me what he means please :)

OpenStudy (anonymous):

answer for b) the two balls have the same velocity = 24.5 m\s

OpenStudy (anonymous):

basically solving for x(t)=0=.5at^2+v_ot+x_o for t

OpenStudy (anonymous):

for both instances and then finding the difference in time

OpenStudy (anonymous):

time 1 is 4? sorry...i'm so slow...i can't find my scientific calculator...i'm such an idiot..:(

OpenStudy (anonymous):

answer c) is 23.49 m

OpenStudy (anonymous):

how to compute for time 2? \[0=(-14.7)t _{2}-(.5)(9.8)^2+0\] @completeidiot ?

OpenStudy (anonymous):

quadratic equation and x_o is not equal to 0 because you're throw it off a high ledge with certain height also you're missing t_2 next to acceleration

OpenStudy (anonymous):

\[-\frac{1}{2}9.8t_2^2-14.7t_2+19.6=0\]

OpenStudy (anonymous):

oh..ok..

OpenStudy (anonymous):

t_2 = 1? t_1= 4?

OpenStudy (anonymous):

@K0GA am i correct?

OpenStudy (anonymous):

yea u r perfect

OpenStudy (anonymous):

but there are another solution shorter

OpenStudy (anonymous):

because There is no air resistance

OpenStudy (anonymous):

so Difference in time will be Time required up and down to balcony

OpenStudy (anonymous):

you got it ?!

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

you do you mean by ok?!

OpenStudy (anonymous):

i got what you're saying...

OpenStudy (anonymous):

if we neglect air resistance

OpenStudy (anonymous):

when i throw the first ball up it will go up until rest then we go down right ?!

OpenStudy (anonymous):

when it arrive the same place its velocity will be the same

OpenStudy (anonymous):

right ?!

OpenStudy (anonymous):

yeA

OpenStudy (anonymous):

ok so when it get back to the same place with the same velocity down it will be like the second stone

OpenStudy (anonymous):

so the difference in time will be the time required to go up and down to balcony

OpenStudy (anonymous):

\[v=v0+at\]

OpenStudy (anonymous):

\[-14.7=14.7-9.8t\]

OpenStudy (anonymous):

\[9.8t=29.4\]

OpenStudy (anonymous):

wat's the answer to that? sorry...just lost my calcu...

OpenStudy (anonymous):

\[t=3 \sec\]

OpenStudy (anonymous):

a) 3 sec

OpenStudy (anonymous):

is that your answer ?!

OpenStudy (anonymous):

wait...from idiot's solution...i got 4 :(

OpenStudy (anonymous):

you got 4 sec for first stone ?!

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

and 1 sec for first stone . right !

OpenStudy (anonymous):

so time difference is 3 sec

OpenStudy (anonymous):

ahh...you're good! :)

OpenStudy (anonymous):

you understand what i said ?

OpenStudy (anonymous):

yes...but is that solution enough? for letter a?

OpenStudy (anonymous):